OpenStudy (anonymous):

Factor completely: 3x – 12

OpenStudy (hoblos):

= 3(x-4)

OpenStudy (anonymous):

3x – 12 common number is 3 3x – 12= 3(x-4)

OpenStudy (anonymous):

Common factor is more proper right?